# SAGA: Source Attribution of Generative AI Videos

### v1

@@ 5 Conclusion
We introduced SAGA, the first comprehensive framework designed for the critical task of multi-granular source attribution of AI-generated videos, moving beyond inadequate binary detection. By combining a novel video transformer with features from a vision foundation model and a data-efficient two-stage contrastive training strategy, SAGA achieves state-of-the-art performance across five attribution levels—from binary to fine-grained generator ID—even with only 0.5% labeled data and in cross-dataset setups. Our introduction of Temporal Attention Signatures (T-Sigs) provides novel interpretability, visually explaining why generators are distinguishable. SAGA establishes a robust benchmark for AI video provenance, offering crucial capabilities for digital forensics and the responsible governance of generative AI.

### v2

@@ 5 Conclusion
We introduced SAGA, the first comprehensive framework designed for the critical task of multi-granular source attribution of AI-generated videos, moving beyond inadequate binary detection. By combining a novel video transformer with features from a vision foundation model and a data-efficient two-stage contrastive training strategy, SAGA achieves state-of-the-art performance across five attribution levels, from binary to fine-grained generator ID, even with only 0.5% labeled data and in cross-dataset setups. Our introduction of Temporal Attention Signatures (T-Sigs) provides novel interpretability, visually explaining why generators are distinguishable. SAGA establishes a robust benchmark for AI video provenance, offering crucial capabilities for digital forensics and the responsible governance of generative AI.
Acknowledgements: This work was supported in part by funding from YouTube (Google LLC).